OpenStudy (anonymous):

how do you find the slope-intercept form of the equation of the line that passes through (-5,3) and is parallel to y=4x-7

OpenStudy (anonymous):

First, do you know the slope of your line?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

Wait, how did you come up with "-5"?

OpenStudy (anonymous):

no......the slope of a line in y = mx + b form, is in the m position. so look at the equation, y = 4x - 7.....what is the slope ?

OpenStudy (anonymous):

Yeah, Kellie's right. Do you know "Point-Slope Form"? Given a slope "m" and a point P \[y - y _{0} = m(x - x _{0})\]

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.